The fundamental equations in an economy are given as:

Consumption function C =200+0.8yd

Investment function. I=300

Tax. T=120

Government expenditure. G=200

Exports. X=100

Imports M=0.05y

Find the following.

1.The equilibrium level of income.

2.The net exports.





1
Expert's answer
2021-05-06T13:23:55-0400

We have given that,

"C = 200+0.8Y_d"

"I=300"

"T=120"

"G=200"

"X=100"

"M=0.05Y"

Consumption Function is C = 200+0.8Y where Y in the income in the economy

a.) At equilibrium level of income

"Y= AE(Aggregate \hspace{1mm}Expenditure)----(1)"

"AE = C + I + G + X \u2013 M"

"AE = 200+0.8Y_d + 300 +200 - 0.05Y"

"AE = 200+ 0.8(Y-T)+300+200-0.05Y"

"AE = 200+0.8(Y-120) +300+200-0.05Y"

Using equation(1),

"Y = 200+0.8(Y-120) +300+200-0.05Y"

"Y-0.8Y+0.05Y = 200-96+500"

"0.25Y = 604"

"Y = 2416"


b.) Imports M "= 0.05Y = 120.8"

Net Exports :

"X-M = 100-120.8 = -20.8"

There is deficit in balance trade.